Circular No. 7028

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

AX J0105–722 J. Yokogawa and K. Koyama, Kyoto University, communicate: “An

ASCA observation made on 1996 May 21–23 detected pulsations with a period of 3.34300 ± 0.00003 s from a source at α = 1^{h}05^{m}06^{s}, δ = −72^{o}11′08′′ (equinox 2000.0; ± 1′ at 90-percent confidence), designated as AX J0105– 722. The position of this source coincides with RX J0105.3–7210 and the radio supernova remnant DEM S128 (Filipovic et al. 1998, A.Ap. Suppl. 127, 119). However, the x-ray spectrum with a photon index of 2.2 ± 0.3 is harder than is usually seen in supernova remnants. Since the significance of the pulse detection is 99.5 percent, we encourage independent confirmations of the pulsation.”

SUPERNOVAE L. Germany reports, on behalf of the the Mount Stromlo Abell Cluster

Supernova Search Team (cf. IAUC 6639), types of two supernovae and several galaxy redshifts obtained from spectra (resolution 0.4 nm) taken with the Mount Stromlo 2.3-m telescope on Oct. 10 and 11. The following redshifts were obtained for the host galaxies of the corresponding supernovae: 1998do, 0.092; 1998dp, 0.057; 1998du, 0.076; 1998dv, 0.156; 1998dw, 0.050. SN 1998dz has no visible host to magnitude R = 21 and is a type-Ia event, 30 days past maximum light, at z ∼ 0.09. SN 1998ea is a type-Ia supernova at z = 0.057, 20 days past maximum light.

NOVA IN M31 N. N. Samus, Sternberg Astronomical Institute, reports the discovery

by R. Inasaridze, G. Kimeridze, and M. Nikolashvili at the Abastumani Observatory of an apparent nova in M31 located at α = 0^{h}42^{m}33^{s}.93, δ = +41^{o}04′56′′.8 (equinox 2000.0). The new star was at photographic magnitudes 18.3 on Aug. 26 and 16.6 on Aug. 29; it has been subsequently confirmed by A. Alksnis (Latvian Radioastrophysical Observatory) and A. Sharov (Sternberg Institute) on Baldone and Crimean plates. Sharov also confirms that the object is not present on a number of earlier photographs or in Hodge’s Atlas of M31.

COMET 21P/GIACOBINI-ZINNER Visual m_{1} estimates by J. Bortle, Stormville, NY (0.41-m refl.): Aug.

28.06 UT, 11.6; Sept. 14.03, 10.8; 18.03, 10.4; 29.02, 9.9; Oct. 12.01, 9.6.

1998 October 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green